The following information has been taken from the books of Lobelia Traders, a general trading shop in Braamfontein, for the month of April 2020: Lobelia Traders is a registered VAT vendor and trades only with registered VAT Vendors. The business charges 15% VAT on all its sales.


1
Expert's answer
2021-04-28T07:51:43-0400

If the VAT payable is considered to be the difference between the outgoing VAT and the incoming VAT, let the income with VAT be 15,000, and the expenses with VAT 10,000.

Outgoing VAT:

"15 000-\frac{15 00}{1.15}=1956.52"

Incoming VAT:

"10 000-\frac{10 000}{1.15}=1304.35"

"VAT payable=Outgoing VAT-Incoming VAT=1956.53-1304.35=652.17"